# 2.5cm2.5cm2cm2cm

### v1

@@ Abstract.
We use vector bundles to study the locus of totally mixed Nash equilibria of an $n$-player game in normal form, which we call Nash equilibrium scheme. When the payoff tensor format is balanced, we study the Nash discriminant variety, i.e., the algebraic variety of games whose Nash equilibrium scheme is nonreduced or has a positive dimensional component. We prove that this variety has codimension one. We classify all components of the Nash equilibrium scheme of binary three-player games. We prove that if the payoff tensor is of boundary format, then the Nash discriminant variety has two components: an irreducible hypersurface and a larger-codimensional component. A generic game with an unbalanced payoff tensor format does not admit totally mixed Nash equilibria. We define the Nash resultant variety of games admitting a positive number of totally mixed Nash equilibria. We prove that it is irreducible and determine its codimension and degree.
@@ 1. Introduction
[…]
We present specific computations for this variety for $(2,2,2)$ and $(2,2,3)$ games in Example 3.9 and Example 3.10. Section 3.2 is dedicated to the study of the Nash discriminant variety of $(2,2,2)$ games, which is an irreducible hypersurface of degree 6. A singular point of the Nash discriminant variety corresponds to a game whose Nash equilibrium scheme is either a line, a nonsingular conic, or a nonsingular cubic. We further compute all the irreducible components of the singular strata of this Nash discriminant variety. The computations are done with Macaulay2 [GS97], and the details are explained on the Mathrepo repository [APS25] with illustrative images. Additionally, we analyze the Nash discriminant variety for these games in the boundary format (Section 3.3). In this case, the Nash discriminant variety can be written as the union of an irreducible hypersurface (with a known degree) and a component of codimension bigger than or equal to two (Theorem 3.26).
@@ 2.2. Preliminaries
In this section, we define an $n$-player game in normal form and the notion of Nash equilibria. We show how globally generated vector bundles (2.4) over the product of projective spaces ${\mathbb{P}}^{\boldsymbol{d}}$ provide an elegant description of the totally mixed Nash equilibria of an $n$-player game. This allows us to give an alternative proof of [MM97, Theorem 3.3] for the maximal number of totally mixed Nash equilibria of generic games, see Theorem 2.7. Any Nash equilibrium gives rise to a totally mixed Nash equilibrium of a smaller game obtained by eliminating unused strategies. Thus, it is still interesting to consider totally mixed Nash equilibria, as, for example, in the generic case, the maximal number of them gives a lower bound for the number of Nash equilibria of the smaller game. The completely (totally) mixed games, i.e., the games with only totally mixed Nash equilibria, are also an interest of study for game theorists (e.g., [Bub79], [CPR74]).
@@
For a given positive integer $n\geq 2$, let $[n]$ be the set of players. If ${\boldsymbol{d}}=(d_{1},\ldots,d_{n})$ with $d_{i}\geq 2$, then we interpret $[d_{i}]$ as the set of strategies for player $i$, and let $I=\prod_{i=1}^{n}[d_{i}]$. For the selected strategies ${\boldsymbol{j}}=(j_{1},\ldots,j_{n})\in I$ of the players, we write $x_{{\boldsymbol{j}}}^{(i)}$ for the payoff for player $i$, and let $X^{(i)}$ be the payoff tensor $(x_{{\boldsymbol{j}}}^{(i)})_{{\boldsymbol{j}}\in I}$. If $X$ denotes the collection $(X^{(1)},\ldots,X^{(n)})\in V^{\oplus n}$ of payoff tensors, then the triplet $([n],I,X)$ is called an $n$-player game in normal form of format ${\boldsymbol{d}}$. For the sake of simplicity, we use $X$ to represent the game $([n],I,X)$.
@@
This paper deals with games of arbitrary numbers of players. However, most examples are concerned with three-player games. Thus, we discuss how we represent a tensor of order $3$.
@@
Figure 1. $3$-player game of format $(3,3,3)$
@@
$\Delta_{i}\coloneqq\left\{\pi^{(i)}\in{\mathbb{R}}^{d_{i}}\,\left|\,\mbox{$\sum_{j=1}^{d_{i}}\pi_{j}^{(i)}=1$, $\pi_{j}^{(i)}\geq 0$ for all $j$}\right.\right\}\,.$
@@
the $2\times 2$ minors of $[X^{(i)}({\boldsymbol{\pi}}_{-i})\mid{\boldsymbol{1}}]$ are
@@
$\begin{cases}\varDelta f_{1,k}^{(i)}=0&\text{for all $i\in[n]$ and $k\in[d_{i}]\setminus\{1\}$,}\\ \pi_{k}^{(i)}>0&\text{for all $i\in[n]$ and $k\in[d_{i}]$,}\\ \sum_{k=1}^{d_{i}}\pi_{k}^{(i)}=1&\text{for all $i\in[n]$.}\end{cases}$ (2.3)
@@
$\begin{matrix}[l]f_{1}^{(1)}&=&-20\,\pi_{1}^{(2)}\pi_{1}^{(3)}-4\,\pi_{2}^{(2)}\pi_{1}^{(3)}-12\,\pi_{3}^{(2)}\pi_{1}^{(3)}-16\,\pi_{1}^{(2)}\pi_{2}^{(3)}+8\,\pi_{2}^{(2)}\pi_{2}^{(3)}+4\,\pi_{3}^{(2)}\pi_{2}^{(3)}+12\,\pi_{1}^{(2)}\pi_{3}^{(3)}+8\,\pi_{2}^{(2)}\pi_{3}^{(3)}-4\,\pi_{3}^{(2)}\pi_{3}^{(3)}\\ f_{2}^{(1)}&=&8\,\pi_{1}^{(2)}\pi_{1}^{(3)}+20\,\pi_{2}^{(2)}\pi_{1}^{(3)}+20\,\pi_{3}^{(2)}\pi_{1}^{(3)}-20\,\pi_{1}^{(2)}\pi_{2}^{(3)}-20\,\pi_{2}^{(2)}\pi_{2}^{(3)}+16\,\pi_{3}^{(2)}\pi_{2}^{(3)}+12\,\pi_{1}^{(2)}\pi_{3}^{(3)}-4\,\pi_{2}^{(2)}\pi_{3}^{(3)}-23\,\pi_{3}^{(2)}\pi_{3}^{(3)}\\ f_{3}^{(1)}&=&-76\,\pi_{1}^{(2)}\pi_{1}^{(3)}+4\,\pi_{2}^{(2)}\pi_{1}^{(3)}-4\,\pi_{3}^{(2)}\pi_{1}^{(3)}+8\,\pi_{1}^{(2)}\pi_{2}^{(3)}+4\,\pi_{2}^{(2)}\pi_{2}^{(3)}-4\,\pi_{3}^{(2)}\pi_{2}^{(3)}+12\,\pi_{1}^{(2)}\pi_{3}^{(3)}-4\,\pi_{2}^{(2)}\pi_{3}^{(3)}+8\,\pi_{3}^{(2)}\pi_{3}^{(3)}\vspace{5pt}\\ f_{1}^{(2)}&=&6\,\pi_{1}^{(1)}\pi_{1}^{(3)}-2\,\pi_{2}^{(1)}\pi_{1}^{(3)}+2\,\pi_{3}^{(1)}\pi_{1}^{(3)}-8\,\pi_{1}^{(1)}\pi_{2}^{(3)}-2\,\pi_{2}^{(1)}\pi_{2}^{(3)}+4\,\pi_{3}^{(1)}\pi_{2}^{(3)}-10\,\pi_{1}^{(1)}\pi_{3}^{(3)}+10\,\pi_{2}^{(1)}\pi_{3}^{(3)}-2\,\pi_{3}^{(1)}\pi_{3}^{(3)}\\ f_{2}^{(2)}&=&-2\,\pi_{1}^{(1)}\pi_{1}^{(3)}-10\,\pi_{3}^{(1)}\pi_{1}^{(3)}+4\,\pi_{1}^{(1)}\pi_{2}^{(3)}+6\,\pi_{2}^{(1)}\pi_{2}^{(3)}-8\,\pi_{3}^{(1)}\pi_{2}^{(3)}-2\,\pi_{1}^{(1)}\pi_{3}^{(3)}+6\,\pi_{2}^{(1)}\pi_{3}^{(3)}-5\,\pi_{3}^{(1)}\pi_{3}^{(3)}\\ f_{3}^{(2)}&=&4\,\pi_{1}^{(1)}\pi_{1}^{(3)}-10\,\pi_{2}^{(1)}\pi_{1}^{(3)}+4\,\pi_{3}^{(1)}\pi_{1}^{(3)}-6\,\pi_{1}^{(1)}\pi_{2}^{(3)}+4\,\pi_{2}^{(1)}\pi_{2}^{(3)}-2\,\pi_{3}^{(1)}\pi_{2}^{(3)}-2\,\pi_{2}^{(1)}\pi_{3}^{(3)}+2\,\pi_{3}^{(1)}\pi_{3}^{(3)}\vspace{5pt}\\ f_{1}^{(3)}&=&-8\,\pi_{1}^{(1)}\pi_{1}^{(2)}-6\,\pi_{2}^{(1)}\pi_{1}^{(2)}+4\,\pi_{3}^{(1)}\pi_{1}^{(2)}-4\,\pi_{1}^{(1)}\pi_{2}^{(2)}-10\,\pi_{2}^{(1)}\pi_{2}^{(2)}+10\,\pi_{3}^{(1)}\pi_{2}^{(2)}+6\,\pi_{1}^{(1)}\pi_{3}^{(2)}+6\,\pi_{2}^{(1)}\pi_{3}^{(2)}-2\,\pi_{3}^{(1)}\pi_{3}^{(2)}\\ f_{2}^{(3)}&=&-8\,\pi_{1}^{(1)}\pi_{1}^{(2)}-4\,\pi_{3}^{(1)}\pi_{1}^{(2)}-4\,\pi_{1}^{(1)}\pi_{2}^{(2)}-10\,\pi_{2}^{(1)}\pi_{2}^{(2)}+2\,\pi_{2}^{(1)}\pi_{3}^{(2)}+14\,\pi_{3}^{(1)}\pi_{3}^{(2)}\\ f_{3}^{(3)}&=&-10\,\pi_{1}^{(1)}\pi_{1}^{(2)}+2\,\pi_{2}^{(1)}\pi_{1}^{(2)}+10\,\pi_{1}^{(1)}\pi_{2}^{(2)}+4\,\pi_{2}^{(1)}\pi_{2}^{(2)}-6\,\pi_{3}^{(1)}\pi_{2}^{(2)}+10\,\pi_{1}^{(1)}\pi_{3}^{(2)}-10\,\pi_{2}^{(1)}\pi_{3}^{(2)}+3\,\pi_{3}^{(1)}\pi_{3}^{(2)}\,.\end{matrix}$
@@
It follows immediately from Definitions 2.3 and 2.6 that ${\boldsymbol{\pi}}=(\pi^{(1)},\ldots,\pi^{(n)})\in\prod_{i=1}^{d_{i}}\Delta_{d_{i}-1}^{\circ}$ is a totally mixed Nash equilibrium of the game $X=(X^{(1)},\ldots,X^{(n)})$ if the corresponding point $[{\boldsymbol{\pi}}]\coloneqq([\pi^{(1)}],\ldots,[\pi^{(n)}])$ of ${\mathbb{P}}^{\boldsymbol{d}}$ lies in the Nash equilibrium scheme ${\mathcal{Z}}_{X}$ of $X$.
@@ 2.3. Vector bundles
If $X=(X^{(1)},\ldots,X^{(n)})\in V^{\oplus n}$, then the $2\times 2$ minors $\varDelta f_{k,\ell}^{(i)}$ of the matrix $[X^{(i)}(\pi_{-i})\mid{\boldsymbol{1}}]$ are multihomogeneous of multidegree ${\boldsymbol{1}}_{i}$. Thus, we view them as elements of the cohomology group $H^{0}({\mathbb{P}}^{\boldsymbol{d}},{\mathcal{O}}({\boldsymbol{1}}_{i}))$ of the line bundle ${\mathcal{O}}({\boldsymbol{1}}_{i})$ on ${\mathbb{P}}^{\boldsymbol{d}}$ and $r_{X}^{(i)}\coloneqq(\varDelta f_{1,2}^{(i)},\ldots,\varDelta f_{1,d_{i}}^{(i)})$ as an element of the cohomology group $H^{0}({\mathbb{P}}^{\boldsymbol{d}},{\mathcal{O}}({\boldsymbol{1}}_{i})^{\oplus(d_{i}-1)})$ of the direct sum ${\mathcal{O}}({\boldsymbol{1}}_{i})^{\oplus(d_{i}-1)}$ of $d_{i}-1$ copies of ${\mathcal{O}}({\boldsymbol{1}}_{i})$. Therefore, $r_{X}\coloneqq(r_{X}^{(1)},\ldots,r_{X}^{(n)})$ can be thought of as a global section of the following vector bundle of rank $D\coloneqq\sum_{i=1}^{n}(d_{i}-1)$ on ${\mathbb{P}}^{\boldsymbol{d}}$:
@@
Since the map from $V^{\oplus n}$ to $H^{0}({\mathbb{P}}^{\boldsymbol{d}},E)$ defined by sending $X$ to $r_{X}$ is onto and since $E$ is globally generated, there exist $X\in V^{\oplus n}$ whose Nash equilibrium schemes are nonsingular of the expected codimensions, and such $X$ form an open subset of $V^{\oplus n}$ (the complement of which will be studied in Section 3). If $X$ is “generic” (i.e., if $X$ belongs to such an open subset) and if $\int_{{\mathbb{P}}^{\boldsymbol{d}}}c_{D}(E)$ denotes the degree of the top Chern class $c_{D}(E)$ of $E$, then $\int_{{\mathbb{P}}^{\boldsymbol{d}}}c_{D}(E)\geq 0$ [EH16, Section 5.5.1], and this is an equality if and only if ${\mathcal{Z}}_{X}=\emptyset$. We use this description to give an alternative formulation and proof to the following well-known result on the maximal number of totally mixed Nash equilibria of a generic game:
@@
Furthermore, if ${\mathcal{Z}}_{X}\neq\emptyset$, then ${\mathcal{Z}}_{X}$ has dimension 00, and its degree is equal to $c({\boldsymbol{d}})$. In particular, the number of distinct totally mixed Nash equilibria of $X$ is bounded above by $c({\boldsymbol{d}})$.
@@
The second and third assertions are immediate consequences of (1), (2), and (3). Thus, we focus on the proof of the equivalence among (1), (2), and (3).
If $c_{t}({\mathcal{O}}({\boldsymbol{1}}_{i}))$ denotes the Chern polynomial of ${\mathcal{O}}({\boldsymbol{1}}_{i})$, then $c_{t}({\mathcal{O}}({\boldsymbol{1}}_{i}))=1+\hat{h}_{i}t$, and hence $c_{t}({\mathcal{O}}({\boldsymbol{1}}_{i})^{\oplus(d_{i}-1)})=(1+\hat{h}_{i}t)^{d_{i}-1}$. So, we have $c_{t}(E)=\prod_{i=1}^{n}(1+\hat{h}_{i}t)^{d_{i}-1}$. Since the $D$th Chern class of $E$ is the coefficient of $t^{D}$ in $c_{t}(E)$, the degree $\int_{{\mathbb{P}}^{\boldsymbol{d}}}c_{D}(E)$ of $c_{D}(E)$ is equal to the coefficient $c({\boldsymbol{d}})$ of $\prod_{i=1}^{n}h_{i}^{d_{i}-1}$ in $\prod_{i=1}^{n}\hat{h}_{i}$. This proves the equivalence between (1) and (2), and hence it remains to show the equivalence between (2) and (3).
@@
If each player of an $n$-player game has $2$ strategies, then

### v2

# A vector bundle approach to Nash equilibria

@@ Abstract.
We use vector bundles to study the locus of totally mixed Nash equilibria of an $n$-player game in normal form, which we call the Nash equilibrium scheme. When the payoff tensor format is balanced, we study the Nash discriminant variety, i.e., the algebraic variety of games whose Nash equilibrium scheme is nonreduced or has a positive dimensional component. We prove that this variety has codimension one. We classify all possible components of the Nash equilibrium scheme for a binary three-player game. We prove that if the payoff tensor is of boundary format, then the Nash discriminant variety has two components: an irreducible hypersurface and a larger-codimensional component. A generic game with an unbalanced payoff tensor format does not admit totally mixed Nash equilibria. We define the Nash resultant variety of games admitting a positive number of totally mixed Nash equilibria. We prove that it is irreducible and determine its codimension and degree.
@@ 1. Introduction
For ${\boldsymbol{d}}=(d,d)$ games, $\Delta({\boldsymbol{d}})$ has two irreducible components of codimension two and degree $\binom{d}{2}$. Otherwise, the real part of the Nash discriminant variety $\Delta({\boldsymbol{d}})$ has codimension one.
We present specific computations for this variety for $(2,2,2)$ and $(2,2,3)$ games in Example 3.9 and Example 3.10. Section 3.2 is dedicated to the study of the Nash discriminant variety of $(2,2,2)$ games.
Theorem (Theorem 3.11, Proposition 3.15).
The Nash discriminant variety $\Delta({\boldsymbol{d}})$ for ${\boldsymbol{d}}=(2,2,2)$ is an irreducible hypersurface of degree 6 in $\mathbb{P}^{11}$. A nonsingular point of $\Delta({\boldsymbol{d}})$ corresponds to a game $X$ whose Nash equilibrium scheme ${\mathcal{Z}}_{X}$ is a nonreduced point of multiplicity two. A game $X$ lies in the singular locus of $\Delta({\boldsymbol{d}})$ if and only if its Nash equilibrium scheme ${\mathcal{Z}}_{X}$ contains a line, a nonsingular conic, or a nonsingular cubic.
We further compute all the irreducible components of the singular strata of this Nash discriminant variety $\Delta(2,2,2)$. The computations are done with Macaulay2 [GS97], and the details are explained on the Mathrepo repository [APS25] with illustrative images. A summary of these results can also be found in Table 2 and Table 4. Additionally, we analyze the Nash discriminant variety for these games in the boundary format (Section 3.3). In this case, the Nash discriminant variety can be written as the union of an irreducible hypersurface (with a known degree) and a component of codimension bigger than or equal to two (Theorem 3.26).
@@ 2.2. Preliminaries
In this section, we define an $n$-player game in normal form and the notion of Nash equilibria. We show how globally generated vector bundles (2.4) over the product of projective spaces ${\mathbb{P}}^{\boldsymbol{d}}$ provide an elegant description of the totally mixed Nash equilibria of an $n$-player game. This allows us to give an alternative proof of [MM97, Theorem 3.3] for the maximal number of totally mixed Nash equilibria of generic games, see Theorem 2.8. Any Nash equilibrium gives rise to a totally mixed Nash equilibrium of a smaller game obtained by eliminating unused strategies. Thus, it is still interesting to consider totally mixed Nash equilibria, as, for example, in the generic case, the maximal number of them gives a lower bound for the number of Nash equilibria of the smaller game. The completely (totally) mixed games, i.e., the games with only totally mixed Nash equilibria, are also of interest to game theorists (e.g., [Bub79], [CPR74]).
@@
For a given positive integer $n\geq 2$, let $[n]$ be the set of players. If ${\boldsymbol{d}}=(d_{1},\ldots,d_{n})$ with $d_{i}\geq 2$, then we interpret $d_{i}]$ as the set of strategies for player $i$, and let $I=\prod_{i=1}^{n}[d_{i}]$. For the selected strategies ${\boldsymbol{j}}=(j_{1},\ldots,j_{n})\in I$ of the players, we write $x_{{\boldsymbol{j}}}^{(i)}$ for the payoff for player $i$, and let $X^{(i)}$ be the payoff tensor $(x_{{\boldsymbol{j}}}^{(i)})_{{\boldsymbol{j}}\in I}$. If $X$ denotes the collection $(X^{(1)},\ldots,X^{(n)})\in V^{\oplus n}$ of payoff tensors, then the triplet $([n],I,X)$ is called an $n$-player game in normal form of format $\boldsymbol{d}$}. For the sake of simplicity, we use $X$ to represent the game $([n],I,X)$.
@@
This paper addresses games with an arbitrary number of players. However, most examples concern three-player games. Thus, we discuss how we represent a tensor of order $3$.
@@
Figure 1. 33-player game of format $(3,3,3)$
@@
[…]
@@
the $2\times 2$ minors of $X^{(i)}({\boldsymbol{\pi}}_{-i})\mid{\boldsymbol{1}}]$ are
@@
[…]
@@
$\begin{matrix}[l]f_{1}^{(1)}&=&-20\,\pi_{1}^{(2)}\pi_{1}^{(3)}-4\,\pi_{2}^{(2)}\pi_{1}^{(3)}-12\,\pi_{3}^{(2)}\pi_{1}^{(3)}-16\,\pi_{1}^{(2)}\pi_{2}^{(3)}+8\,\pi_{2}^{(2)}\pi_{2}^{(3)}+4\,\pi_{3}^{(2)}\pi_{2}^{(3)}+12\,\pi_{1}^{(2)}\pi_{3}^{(3)}+8\,\pi_{2}^{(2)}\pi_{3}^{(3)}-4\,\pi_{3}^{(2)}\pi_{3}^{(3)}\\ f_{2}^{(1)}&=&8\,\pi_{1}^{(2)}\pi_{1}^{(3)}+20\,\pi_{2}^{(2)}\pi_{1}^{(3)}+20\,\pi_{3}^{(2)}\pi_{1}^{(3)}-20\,\pi_{1}^{(2)}\pi_{2}^{(3)}-20\,\pi_{2}^{(2)}\pi_{2}^{(3)}+16\,\pi_{3}^{(2)}\pi_{2}^{(3)}+12\,\pi_{1}^{(2)}\pi_{3}^{(3)}-4\,\pi_{2}^{(2)}\pi_{3}^{(3)}-23\,\pi_{3}^{(2)}\pi_{3}^{(3)}\\ f_{3}^{(1)}&=&-76\,\pi_{1}^{(2)}\pi_{1}^{(3)}+4\,\pi_{2}^{(2)}\pi_{1}^{(3)}-4\,\pi_{3}^{(2)}\pi_{1}^{(3)}+8\,\pi_{1}^{(2)}\pi_{2}^{(3)}+4\,\pi_{2}^{(2)}\pi_{2}^{(3)}-4\,\pi_{3}^{(2)}\pi_{2}^{(3)}+12\,\pi_{1}^{(2)}\pi_{3}^{(3)}-4\,\pi_{2}^{(2)}\pi_{3}^{(3)}+8\,\pi_{3}^{(2)}\pi_{3}^{(3)}\vskip 5.0pt\\ f_{1}^{(2)}&=&6\,\pi_{1}^{(1)}\pi_{1}^{(3)}-2\,\pi_{2}^{(1)}\pi_{1}^{(3)}+2\,\pi_{3}^{(1)}\pi_{1}^{(3)}-8\,\pi_{1}^{(1)}\pi_{2}^{(3)}-2\,\pi_{2}^{(1)}\pi_{2}^{(3)}+4\,\pi_{3}^{(1)}\pi_{2}^{(3)}-10\,\pi_{1}^{(1)}\pi_{3}^{(3)}+10\,\pi_{2}^{(1)}\pi_{3}^{(3)}-2\,\pi_{3}^{(1)}\pi_{3}^{(3)}\\ f_{2}^{(2)}&=&-2\,\pi_{1}^{(1)}\pi_{1}^{(3)}-10\,\pi_{3}^{(1)}\pi_{1}^{(3)}+4\,\pi_{1}^{(1)}\pi_{2}^{(3)}+6\,\pi_{2}^{(1)}\pi_{2}^{(3)}-8\,\pi_{3}^{(1)}\pi_{2}^{(3)}-2\,\pi_{1}^{(1)}\pi_{3}^{(3)}+6\,\pi_{2}^{(1)}\pi_{3}^{(3)}-5\,\pi_{3}^{(1)}\pi_{3}^{(3)}\\ f_{3}^{(2)}&=&4\,\pi_{1}^{(1)}\pi_{1}^{(3)}-10\,\pi_{2}^{(1)}\pi_{1}^{(3)}+4\,\pi_{3}^{(1)}\pi_{1}^{(3)}-6\,\pi_{1}^{(1)}\pi_{2}^{(3)}+4\,\pi_{2}^{(1)}\pi_{2}^{(3)}-2\,\pi_{3}^{(1)}\pi_{2}^{(3)}-2\,\pi_{2}^{(1)}\pi_{3}^{(3)}+2\,\pi_{3}^{(1)}\pi_{3}^{(3)}\vskip 5.0pt\\ f_{1}^{(3)}&=&-8\,\pi_{1}^{(1)}\pi_{1}^{(2)}-6\,\pi_{2}^{(1)}\pi_{1}^{(2)}+4\,\pi_{3}^{(1)}\pi_{1}^{(2)}-4\,\pi_{1}^{(1)}\pi_{2}^{(2)}-10\,\pi_{2}^{(1)}\pi_{2}^{(2)}+10\,\pi_{3}^{(1)}\pi_{2}^{(2)}+6\,\pi_{1}^{(1)}\pi_{3}^{(2)}+6\,\pi_{2}^{(1)}\pi_{3}^{(2)}-2\,\pi_{3}^{(1)}\pi_{3}^{(2)}\\ f_{2}^{(3)}&=&-8\,\pi_{1}^{(1)}\pi_{1}^{(2)}-4\,\pi_{3}^{(1)}\pi_{1}^{(2)}-4\,\pi_{1}^{(1)}\pi_{2}^{(2)}-10\,\pi_{2}^{(1)}\pi_{2}^{(2)}+2\,\pi_{2}^{(1)}\pi_{3}^{(2)}+14\,\pi_{3}^{(1)}\pi_{3}^{(2)}\\ f_{3}^{(3)}&=&-10\,\pi_{1}^{(1)}\pi_{1}^{(2)}+2\,\pi_{2}^{(1)}\pi_{1}^{(2)}+10\,\pi_{1}^{(1)}\pi_{2}^{(2)}+4\,\pi_{2}^{(1)}\pi_{2}^{(2)}-6\,\pi_{3}^{(1)}\pi_{2}^{(2)}+10\,\pi_{1}^{(1)}\pi_{3}^{(2)}-10\,\pi_{2}^{(1)}\pi_{3}^{(2)}+3\,\pi_{3}^{(1)}\pi_{3}^{(2)}\,.\end{matrix}$
@@
It follows immediately from Definitions 2.3 and 2.6 that ${\boldsymbol{\pi}}=(\pi^{(1)},\ldots,\pi^{(n)})\in\prod_{i=1}^{d_{i}}\Delta_{d_{i}-1}^{\circ}$ is a totally mixed Nash equilibrium of the game $X=(X^{(1)},\ldots,X^{(n)})$ if the corresponding point ${\boldsymbol{\pi}}]\coloneqq([\pi^{(1)}],\ldots,[\pi^{(n)}])$ of ${\mathbb{P}}^{\boldsymbol{d}}$ lies in the Nash equilibrium scheme ${\mathcal{Z}}_{X}$ of $X$.
@@ 2.3. Vector bundles
If $X=(X^{(1)},\ldots,X^{(n)})\in V^{\oplus n}$, then the $2\times 2$ minors $\varDelta f_{k,\ell}^{(i)}$ of the matrix $X^{(i)}(\pi_{-i})\mid{\boldsymbol{1}}]$ are multihomogeneous of multidegree ${\boldsymbol{1}}_{i}$. Thus, we view them as elements of the cohomology group $H^{0}({\mathbb{P}}^{\boldsymbol{d}},{\mathcal{O}}({\boldsymbol{1}}_{i}))$ of the line bundle ${\mathcal{O}}({\boldsymbol{1}}_{i})$ on ${\mathbb{P}}^{\boldsymbol{d}}$ and $r_{X}^{(i)}\coloneqq(\varDelta f_{1,2}^{(i)},\ldots,\varDelta f_{1,d_{i}}^{(i)})$ as an element of the cohomology group $H^{0}({\mathbb{P}}^{\boldsymbol{d}},{\mathcal{O}}({\boldsymbol{1}}_{i})^{\oplus(d_{i}-1)})$ of the direct sum ${\mathcal{O}}({\boldsymbol{1}}_{i})^{\oplus(d_{i}-1)}$ of $d_{i}-1$ copies of ${\mathcal{O}}({\boldsymbol{1}}_{i})$. Therefore, $r_{X}\coloneqq(r_{X}^{(1)},\ldots,r_{X}^{(n)})$ can be thought of as a global section of the following vector bundle of rank $D\coloneqq\sum_{i=1}^{n}(d_{i}-1)$ on ${\mathbb{P}}^{\boldsymbol{d}}$:
@@
The vector bundle $E$ is globally generated, as so are its direct summands. Thus, the degree $\int_{{\mathbb{P}}^{{\boldsymbol{d}}}}c_{D}(E)$ of the top Chern class $c_{D}(E)$ of $E$ is nonnegative [Gee02, Proposition 10]. Moreover, the global sections $f$ of $E$ whose zero schemes ${\mathcal{Z}}(f)$ are nonsingular of the expected codimension (meaning that it is of codimension $D+1$ if $\int_{{\mathbb{P}}^{{\boldsymbol{d}}}}c_{D}(E)=0$; it is of codimension $D$ otherwise) form an open subset $U(E)$ of $H^{0}({\mathbb{P}}^{{\boldsymbol{d}}},E)$ [Ein82, Lemma 2.5].
Proposition 2.7.
The map $\Phi\colon V^{\oplus n}\to H^{0}({\mathbb{P}}^{{\boldsymbol{d}}},E)$ defined by $\Phi(X)=r_{X}$ is an onto linear transformation. In particular, for each $X\in V^{\oplus n}$ there exists a global section $f$ of the vector bundle $E$ of rank $D$ such that ${\mathcal{Z}}_{X}={\mathcal{Z}}(f)$. Furthermore, the subset of elements of $V^{\oplus n}$ whose Nash equilibrium scheme is nonsingular of the expected codimension forms an open subset.
Proof.
For each $i\in\{1,\ldots,n\}$, let $V_{i}$ denote the $i$th direct summand of $V^{\oplus n}$ and let $\Phi_{i}\colon V_{i}\to H^{0}({\mathbb{P}}^{{\boldsymbol{d}}},{\mathcal{O}}({\boldsymbol{1}}_{i})^{\oplus(d_{i}-1)})$ be the map defined by $\Phi_{i}(X^{(i)})=r_{X}^{(i)}=(\varDelta f_{1,2}^{(i)},\ldots,\varDelta f_{1,d_{i}}^{(i)})$. The linearity of $\Phi_{i}$ follows from that of the coefficient $x_{(1,{\boldsymbol{j}}_{-i})}^{(i)}-x_{(\ell,{\boldsymbol{j}}_{-i})}^{(i)}$ of $\pi_{{\boldsymbol{j}}_{-i}}$ in $\varDelta f_{1,\ell}^{(i)}$. These coefficients are linearly independent, which implies the surjectivity of $\Phi_{i}$. Note that $\Phi$ is the direct sum of $\Phi_{1},\ldots,\Phi_{n}$. Thus, it is an onto linear transformation. Therefore, for each $X^{(i)}\in V_{i}$, there exists a global section $f^{(i)}$ of ${\mathcal{O}}({\boldsymbol{1}}_{i})^{\oplus(d_{i}-1)}$ such that $r_{X}^{(i)}=f^{(i)}$. Since $r_{X}^{(i)}$ generated the ideal $J_{i}$ and since $J=\sum_{i=1}^{n}J_{i}$ defines the Nash equilibrium scheme ${\mathcal{Z}}_{X}$ of $X$, the zero scheme of $f=(f^{(1)},\ldots,f^{(n)})\in H^{0}({\mathbb{P}}^{{\boldsymbol{d}}},E)$ coincides with ${\mathcal{Z}}_{X}$.
The subset of elements of $V^{\oplus n}$ whose Nash equilibrium schemes are nonsingular of the expected codimension is the inverse image of $U(E)$ under $\Phi$. Therefore, it is open in $V^{\oplus n}$. ∎
We say that a game $X\in V^{\oplus n}$ is generic if $X$ belongs to the open subset of $V^{\oplus n}$ defined in Proposition 2.7. As formalized in the upcoming Theorem 2.8, there is a dichotomy in the previous notion of genericity, which depends only on the format $\boldsymbol{d}$}. We summarize this in Table 1.
Within boundary format Beyond boundary format
Formulation $d_{n}-1\leq\sum_{i=1}^{n-1}(d_{i}-1)$ $d_{n}-1>\sum_{i=1}^{n-1}(d_{i}-1)$
Generic $X$ $\dim{\mathcal{Z}}_{X}=0$ and ${\mathcal{Z}}_{X}$ reduced of degree $c({\boldsymbol{d}})$ ${\mathcal{Z}}_{X}=\emptyset$
Nongeneric $X$ Nash discriminant variety, Definition 3.2 Nash resultant variety, Definition 3.29
[…]
Table 1. The dictionary for format separation, the generic behavior of the Nash equilibrium scheme ${\mathcal{Z}}_{X}$, and nongeneric behavior of games, depending on the format ${\boldsymbol{d}}=(d_{1},\ldots,d_{n})\in{\mathbb{Z}}_{{\geq 2}}^{n}$ with $d_{1}\leq\cdots\leq d_{n}$.
We use the above description and Proposition 2.7 to give an alternative formulation and proof to the well-known result on the maximal number of totally mixed Nash equilibria of a generic game in terms of the degree of the top Chern class of the vector bundle $E$. Indeed, Mckelvey and McLennan [MM97] compute this maximal number as the mixed volume of the Newton polytopes associated to the generators of the multihomogeneous ideal $J$ defining the Nash equilibrium scheme ${\mathcal{Z}}_{X}$ (Definition 2.6). Equivalently, this number can be computed combinatorially via so-called block derangements [Stu02, Theorem 6.8].
@@
Furthermore, if ${\mathcal{Z}}_{X}\neq\emptyset$, then ${\mathcal{Z}}_{X}$ has dimension 0, and its degree is equal to $c({\boldsymbol{d}})$. In particular, the number of distinct totally mixed Nash equilibria of $X$ is bounded above by $c({\boldsymbol{d}})$.
@@
The second and third assertions are immediate consequences of (1), (2), and (3). Thus, we focus on proving the equivalence of (1), (2), and (3).
If $c_{t}({\mathcal{O}}({\boldsymbol{1}}_{i}))$ denotes the Chern polynomial of ${\mathcal{O}}({\boldsymbol{1}}_{i})$, then $c_{t}({\mathcal{O}}({\boldsymbol{1}}_{i}))=1+\hat{h}_{i}t$, and hence $c_{t}({\mathcal{O}}({\boldsymbol{1}}_{i})^{\oplus(d_{i}-1)})=(1+\hat{h}_{i}t)^{d_{i}-1}$. So, we have $c_{t}(E)=\prod_{i=1}^{n}(1+\hat{h}_{i}t)^{d_{i}-1}$. Since the $D$th Chern class of $E$ is the coefficient of $t^{D}$ in $c_{t}(E)$, the degree $\int_{{\mathbb{P}}^{\boldsymbol{d}}}c_{D}(E)$ of $c_{D}(E)$ is equal to the coefficient $c({\boldsymbol{d}})$ of $\prod_{i=1}^{n}h_{i}^{d_{i}-1}$ in $\prod_{i=1}^{n}\hat{h}_{i}$. This establishes the equivalence between (1) and (2); it remains to show the equivalence between (2) and (3).
@@
If each player of an $n$-player game has 22 strategies, then